OpenStudy (anonymous):

Generate the first 5 terms of this sequence: F(1)=2 and f(2)=3, f(n)=f(n-1)+f(n-2), for n>2. A.2,3,5,7,9 B.2,3,4,7,11 C.2,3,5,9,11 D.2,3,5,8,13 please

OpenStudy (neer2890):

put the value of n=1,2,3,4,5 in f(n)

OpenStudy (neer2890):

for eg., f(n)=f(n-1)+f(n-2) so f(3)=f(3-1)+f(3-2) =f(2)+f(1) =3+2=5